# Triple and binary evolution pathways play equally significant roles in LISA double white dwarf formation

### v1

@@
22institutetext: Radboud Univeristy, Nijmegen 33institutetext: Anton Pannekoek Institute for Astronomy, University of Amsterdam, Science Park 904, 1098, XH, Amsterdam, The Netherlands 44institutetext: Ludwig-Maximilians-Universität München, Geschwister-Scholl-Platz 1, 80539 München, Germany
Triple and binary evolution pathways play equally significant roles in LISA double white dwarf formation
[…]
Self-consistent evolution of a population of Galactic LISA double white dwarfs formed in triples
[…]
@@
[…]
Galactic double white dwarfs will be prominent gravitational-wave sources for the Laser Interferometer Space Antenna (LISA). While previous studies have primarily focused on formation scenarios in which binaries form and evolve in isolation, we present the first detailed study of the role of triple stellar evolution in forming the population of LISA double white dwarfs. In this work, we present the first detailed study of the role of triple stellar evolution in forming the population of LISA double white dwarfs. We use the multiple stellar evolution code (MSE) to model the stellar evolution, binary interactions, and the dynamics of triple star systems then use a Milky Way-like galaxy from the TNG50 simulations to construct a representative sample of LISA double white dwarfs. In our simulations about $7\times 10^{6}$ Galactic double white dwarfs in the LISA frequency bandwidth originate from triple systems, whereas $\sim 4\times 10^{6}$ form from isolated binary stars. The properties of double white dwarfs formed in triples closely resemble those formed from isolated binaries, but we also find a small number of systems $\sim\mathcal{O}(10)$ that reach extreme eccentricities $(>0.9)$, a feature unique to the dynamical formation channels. Our population produces $\approx 10^{4}$ individually resolved double white dwarfs (from triple and binary channels) and an unresolved stochastic foreground below the level of the LISA instrumental noise. About $57\,\%$ of double white dwarfs from triple systems retain a bound third star when entering the LISA frequency bandwidth. However, we expect the tertiary stars to be too distant to have a detectable imprint in the gravitational-wave signal of the inner binary.
@@ 1 Introduction
Observations indicate that triple star systems are common across various stellar evolutionary stages, including main-sequence stars, evolved giant stars, brown dwarfs, and black holes (Kervella et al., 2017; Faherty et al., 2011; Triaud et al., 2020; Lillo-Box et al., 2021; Rivinius et al., 2020). In particular, white dwarfs have been found in triple systems with main-sequence (MS) stars (for a review within 20 pc, see Toonen et al. 2014), other white dwarfs (Maxted et al., 2000; Perpinyà-Vallès et al., 2019), neutron stars (NSs)(Ransom et al., 2014), and brown dwarfs (BDs) (Rebassa-Mansergas et al., 2022). However, the observed number of white dwarfs in triple systems remains significantly lower than theoretical predictions, possibly due to observational biases. Indeed, Shariat et al. (2024) propose that many observed local double white dwarfs could have originated from triples. For example, a recent spectroscopic study of double white dwarfs by Heintz et al. (2024) reveals that the more massive white dwarf companion has a shorter cooling age compared to the less massive one, which contradicts the initial-final mass relation and the assumption that both stars in the binary formed simultaneously from the same molecular cloud. One proposed explanation is that the progenitors of the double white dwarf were originally in a triple system, where the massive white dwarf was formed by the merger of two stars, resulting in a shorter cooling age. Furthermore, triples could significantly contribute to the rate of Type Ia supernovae, making a substantial contribution to that from isolated binary stars (Katz et al., 2011; Hamers et al., 2013; Rajamuthukumar et al., 2023).
@@
Hierarchical triple systems are characterized by a close inner orbit and with a tertiary component in a wider orbit. When the orbits of the inner and outer stars are sufficiently inclined, the gravitational perturbation from the tertiary star can cause large-amplitude von Zeipel-Lidov-Kozai (ZLK) oscillations (von Zeipel, 1909; Lidov, 1962; Kozai, 1962) oscillations of the inner binary eccentricity whilst the semi-major axis is unchanged. This process can play a key role in the formation of close binaries. The combination of ZLK oscillations with dissipative effects like tidal friction (Kiseleva et al., 1998; Eggleton & Kiseleva-Eggleton, 2001; Fabrycky & Tremaine, 2007) and gravitational-wave radiation can lead to a reduction in the inner binary’s orbital semi-major axis. Thus, perturbations from the tertiary star can facilitate close-binary processes such as mass transfer, common-envelope phases, mergers, and collisions in the inner binary (Fabrycky & Tremaine, 2007; Perets & Kratter, 2012; Shappee & Thompson, 2013; Hamers et al., 2013; Michaely & Perets, 2014; Antonini et al., 2017; Toonen et al., 2020; Hamers & Thompson, 2019; Stegmann et al., 2022a, b).
@@
Gravitational-waves from compact double white dwarfs (with frequencies ranging from $10^{-4}$ to $10^{-1}\,\rm Hz$) detectable with the upcoming Laser Interferometer Space Antenna (LISA) mission offer a unique way to explore these triple systems (Amaro-Seoane et al., 2023). By detecting the gravitational-wave signals from double white dwarfs, LISA could uncover a population of systems formed through the triple evolution channel which are inaccessible to electromagnetic observations. This capability has the potential to provide new insights into the formation mechanisms of double white dwarfs, their contribution to Type Ia supernovae (Iben & Tutukov, 1984), and the broader implications for the chemical evolution of the Galaxy (Pagel, 1997).
@@
LISA is expected to detect $\sim\mathcal{O}(10^{6})$ of Galactic double white dwarfs as part of an unresolved confusion gravitational-wave background and individually resolve around $10^{3}$ – $10^{4}$ of the ’loudest’ double white dwarfs (e.g., Korol et al., 2017; Lamberts et al., 2019; Thiele et al., 2023; Wilhelm et al., 2021; Li et al., 2023; Tang et al., 2024). While these previous studies has focussed on double white dwarfs formed from the evolution of isolated binary stars, increasing evidence suggests that hierarchical triple systems, in which a close inner binary is orbited by a distant tertiary companion, may also play a significant role in the formation of double white dwarfs (Toonen et al., 2020; Heintz et al., 2024; Shariat et al., 2024).
@@
There is mounting observational evidence that stars often form with bound companions, with a binary fraction of $30\,\%$ and a triple fraction of $10\,\%$ for F and G-type stars (i.e. with masses $\sim 1\,\mathrm{M_{\odot}}$) (Eggleton & Tokovinin, 2008; Raghavan et al., 2010; Tokovinin, 2014; Moe & Di Stefano, 2017; Offner et al., 2023). Moreover, the inner binaries in triple systems tend to be in closer orbits than the binaries found in isolated systems, which increases the probability for some binary interactions (Toonen et al., 2020). In this paper, we show that triple systems offer a greater number of evolutionary pathways for forming short-period inner binaries compared to isolated binary systems.
@@
Previous studies have explored the potential for detecting tertiary companions in LISA data (e.g., Seto, 2008; Robson et al., 2018; Tamanini & Danielski, 2019). Similar to electromagnetic observations, the motion of the double white dwarf around the center of mass of the triple system modulates the gravitational-wave frequency through the Doppler effect. This modulation causes a periodic shift in the gravitational-wave frequency, oscillating around the intrinsic frequency of the inner binary. Recent studies have focused on leveraging this effect to detect sub-stellar mass tertiaries, such as exoplanets and brown dwarfs (Tamanini & Danielski, 2019; Danielski et al., 2019; Kang et al., 2021; Katz et al., 2022). Thus, previous studies have primarily focused on either the isolated binary population of double white dwarfs or the detection possibilities of the third star. Our work is the first evolutionary population synthesis study of Galactic double white dwarfs resulting from triple evolution. In addition, our study accesses the impact of the triple evolution channel to the LISA’s astrophysical noise background, thereby influencing the $\rho$ of all other gravitational-wave sources.
@@
We aim to quantify the contribution of the triples to the population of double white dwarfs detectable by LISA. We combine population synthesis models using the Multiple Star Evolution (MSE) code (Hamers et al., 2021) with cosmological simulations from the TNG50 project (Nelson et al., 2019; Pillepich et al., 2019) to construct a representative model of the Galactic double white dwarf population. Our study addresses two critical questions: 1) What fraction of double white dwarfs detectable by LISA originates from the triple evolution channel? 2) Can LISA detect the dynamical effects of the third star in these triple systems?
@@
The paper is structured as follows. In Section 2, we explain our methodology. Section 3 describes the evolutionary pathways of triples that lead to the formation of LISA double white dwarfs. We detail the population properties of LISA double white dwarfs from isolated binaries and triples in Section 4, investigating prospects for direct detection of the third star in Section 4.4. Finally, we discuss the results in Section 5 and summarise our findings in Section 6.
@@ 2 Methods
All simulations are performed using the publicly available population synthesis code MSE [^1] https://github.com/hpreece/mse. (Hamers et al., 2021). Our set of simulations of hierarchical triples consists of a main run with a choice of default parameters and three model-variant runs. In each simulated data set, we evolve the triples from the start of the zero-age-main-sequence until a maximum integration time $t_{\rm max}=14\,\rm Gyr$. The main triple data set consists of $10^{5}$ systems, where we adopt a common envelope efficiency parameter $\mathrm{\alpha_{CE}=1}$, and all three stars are assumed to have formed at solar metallicity $Z=\rm Z_{\odot}=0.02$. This population results in $3\times 10^{3}$ LISA double white dwarfs, yielding a $\sim 2\,\%$ Poisson error. In each of the three model-variant runs, we simulate $10^{4}$ systems and either vary the common envelope efficiency parameter as $\mathrm{\alpha_{CE}=0.1}$ and 10 or change the metallicity of the stars to sub-solar $Z=0.1\,\mathrm{Z_{\odot}}$. The effects of the chosen parameters are discussed in Section 5. In addition to the triple runs, we simulate a population of $10^{5}$ isolated binaries to compare the impact of tertiary companions. Additionally, we model the inner binaries of all triples without their tertiary stars to assess their influence on the resulting LISA double white dwarf population. We expect that the isolated binary population differs significantly from the inner binary population of triples. The primary distinction is that the inner binaries have much more compact semi-major axes due to the dynamical stability constraints imposed by the tertiary star.
@@
2.1 Multiple stellar evolution code (MSE)
@@
We use the population synthesis approach using the code MSE to model the stellar evolution, binary interactions (tides, mass transfer, etc.), dynamical perturbations from higher-order companions in multiple systems, and fly-bys from ambient stars. MSE is a C/C++ code with a Python interface that can handle any number of stars as long as they start in a hierarchical arrangement. MSE uses a hybrid approach that switches between the secular approximation for dynamically stable orbits (that satisfy the criterion of Mardling & Aarseth 2001) and direct $N$-body integration using MSTAR (Rantala et al., 2020) for dynamically unstable orbits. Throughout the evolution post-Newtonian terms are included to 2.5 order in the secular approximation and to 3.5 order for the direct $N$-body integration.
@@
To follow the evolution of single stars, MSE relies on the fitting formulae from Hurley et al. (2000); Tout et al. (1996), while binary interactions such as tides, wind mass transfer, stable mass transfer episodes, and common envelope (CE) evolutions are computed using modified prescriptions of Hurley et al. (2002). We briefly explain the physical handling of these key processes below. For more detailed explanations see Hamers et al. (2021).
@@
Stable mass transfer: In MSE, mass transfer stability is determined either by the critical mass ratio criterion or by comparing mass transfer and dynamical timescales. The critical mass ratio depends on the donor’s stellar type (Hamers et al., 2021). We assume fully conservative mass transfer ($\beta_{\mathrm{MT}}=1$), meaning no mass is lost from the system. While MSE generally follows Hurley et al. (2002) for binary interactions, it differs in treating mass transfer in eccentric orbits. Hurley et al. (2002) assumes tides are always efficient in circularizing the orbit. However, in triple systems, the eccentricities are excited secularly. We follow the analytical model from Hamers & Dosopoulou (2019) to model mass transfer at periastron in eccentric orbits.
@@
Common envelope evolution: Unstable mass transfer/CE evolution in MSE follows the $\alpha_{\mathrm{CE}}$ prescription (Paczynski, 1976; van den Heuvel, 1976; Livio & Soker, 1988; Iben & Livio, 1993; Hurley et al., 2002). The code solves for orbital energies before and after the CE phase, parameterizing the envelope ejection efficiency by $\alpha_{\mathrm{CE}}$ and assuming a binding energy factor of $\lambda=1$. For the main runs, we adopt $\alpha_{\mathrm{CE}}=1$ and common envelope time of $10^{3}\,\mathrm{yr}$. The post-CE semi-major axis is determined from the corresponding orbital energy.
@@
Merger/collision: A “failed” CE can result in the merger of two stars. This occurs if the post-CE semi-major axis is too small for either star to avoid Roche lobe overflow. Beyond post-CE mergers, MSE also accounts for physical collisions when the sum of the stellar radii exceeds the semi-major axis or when a periapsis collision occurs in an eccentric orbit. The properties of the merger remnant are assigned following Hurley et al. (2002).
@@
Triple common envelope (TCE) evolution: Mass transfer from a third star onto the inner binary can lead to a TCE. If CE conditions are met for the third star, MSE employs “circumstellar triple CE evolution”, allowing the third star to fill its Roche lobe around the inner binary and undergo unstable mass transfer. This follows a similar approach to the prescription proposed by Comerford & Izzard (2020). The final outer semi-major axis is estimated using an $\alpha_{\mathrm{CE}}$ prescription, assuming the inner binary remains relatively compact. However, CE modeling is inherently uncertain (see Ivanova et al. 2013 for a review), and TCE evolution is even more so, requiring cautious interpretation of results. For detailed hydrodynamical simulations of triple CE outcomes, see e.g. Glanz & Perets (2020).
fly-bys: MSE also includes the gravitational perturbations from stellar fly-bys in the vicinity of the system using the impulsive approximation. The fly-by mass is sampled from a Kroupa initial mass function (Kroupa, 2001), and encounters are randomly sampled within an encounter sphere of radius $R_{\text{enc}}=10^{5}\,au$ with velocities drawn from a Maxwellian velocity distribution with dispersion $\sigma_{\star}=30$ km s-1. The number density of fly-bys accounts for the low-density environments, assuming $n_{\star}=0.1\,\mathrm{pc}^{-3}$. These fly-bys become significant for the evolution of the system if the semi-major axis of the outer orbit exceeds about $10^{3}\,\rm au$ (Jiang & Tremaine, 2010; Grishin & Perets, 2022; Stegmann et al., 2024).

### v2

# The role of triple evolution in the formation of LISA double white dwarfs

@@
22institutetext: Radboud University, Nijmegen, The Netherlands 33institutetext: Anton Pannekoek Institute for Astronomy, University of Amsterdam, Science Park 904, 1098, XH, Amsterdam, The Netherlands 44institutetext: Ludwig-Maximilians-Universität München, Geschwister-Scholl-Platz 1, 80539 München, Germany
@@
Abinaya Swaruba Rajamuthukumar    Valeriya Korol    Jakob Stegmann    Holly Preece    Rüdiger Pakmor    Stephen Justham    Silvia Toonen    Selma E. de Mink
Galactic double white dwarfs will be prominent gravitational wave sources for the Laser Interferometer Space Antenna (LISA). While previous studies have primarily focused on formation scenarios in which binaries form and evolve in isolation, we present the first detailed study of the role of triple stellar evolution in forming the population of LISA double white dwarfs. We used the multiple stellar evolution code (MSE) to model the stellar evolution, binary interactions, and the dynamics of triple star systems and then used a Milky Way-like galaxy from the TNG50 simulations to construct a representative sample of LISA double white dwarfs. In our simulations, about $7\times 10^{6}$ Galactic double white dwarfs in the LISA frequency bandwidth originate from triple systems, whereas $\sim 4\times 10^{6}$ are in isolated binary stars. The properties of double white dwarfs formed in triples closely resemble those formed from isolated binaries, but we also find a small number of systems, $\sim\mathcal{O}(10)$, that reach extreme eccentricities $(>0.9)$, a feature unique to the dynamical formation channels. Our population produces $\sim\mathcal{O}(10^{4})$ individually resolved double white dwarfs (from triple and binary channels) and an unresolved stochastic foreground below the level of the LISA instrumental noise. About $57\,\%$ of the double white dwarfs from triple systems retain a bound third star when entering the LISA frequency bandwidth. However, we expect the tertiary stars to be too distant to have a detectable imprint in the gravitational wave signal of the inner binary.
@@ 1 Introduction
Observations indicate that triple star systems are common across various stellar evolutionary stages, including main-sequence stars, evolved giant stars, brown dwarfs, and black holes (2008MNRAS.389..869E; 2017ApJS..230...15M; 2017A&A...598L...7K; 2020NatAs...4..650T; 2021A&A...653A..40L; 2024Natur.635..316B). In particular, white dwarfs have been found in triple systems with main-sequence stars (for a review within 20 pc, see 2014A&A...562A..14T), other white dwarfs (10.1046/j.1365-8711.2000.03343.x; 2019MNRAS.483..901P), neutron stars (2014Natur.505..520R), and brown dwarfs (2022ApJ...927L..31R). However, the observed number of white dwarfs in triple systems remains significantly lower than theoretical predictions, possibly due to observational biases. Indeed, 2023ApJ...955L..14S; 2024arXiv240706257S propose that many observed local double white dwarfs could have originated from triples. For example, a recent spectroscopic study of wide double white dwarfs by 2024ApJ...969...68H reveals some systems in which the more massive white dwarf companion has a shorter cooling age compared to the less massive one, which contradicts the initial-final mass relation if both stars were formed simultaneously in a non-interacting binary. One proposed explanation is that the progenitors of the double white dwarf were originally in a triple system, where the massive white dwarf was formed by the merger of two stars, resulting in a shorter cooling age. Furthermore, triples could significantly contribute to the rate of Type Ia supernovae, making a substantial contribution to the Type Ia supernova rate from isolated binary stars (2011PhRvL.107r1101K; 2013MNRAS.430.2262H; 2023ApJ...950....9R).
@@
Hierarchical triple systems are characterized by a close inner orbit and with a tertiary component in a wider orbit. When the orbits of the inner and outer stars are sufficiently inclined, the gravitational perturbation from the tertiary star can cause large-amplitude von Zeipel-Lidov-Kozai (ZLK) oscillations (Zeipel1910; 1962P&SS....9..719L; 1962AJ.....67..591K) of the inner binary eccentricity while the semi-major axis remains unchanged. This process can play a key role in the formation of close binaries. The combination of ZLK oscillations with dissipative effects such as tidal friction (10.1046/j.1365-8711.1998.01903.x; Eggleton_2001; 2007ApJ...669.1298F) and gravitational wave radiation can lead to a reduction in the inner binary’s orbital semi-major axis. Thus, perturbations from the tertiary star can facilitate close-binary processes such as mass transfer, common-envelope phases, mergers, and collisions in the inner binary (2007ApJ...669.1298F; 2012ApJ...760...99P; 2013ApJ...766...64S; 2013MNRAS.430.2262H; 2014ApJ...794..122M; 2017ApJ...841...77A; 2020A&A...640A..16T; 2019ApJ...882...24H; 2022MNRAS.516.1406S; 2022PhRvD.106b3014S).
@@
Gravitational waves from compact double white dwarfs (with frequencies ranging from $10^{-4}$ to $10^{-1}\,\rm Hz$) detectable with the upcoming Laser Interferometer Space Antenna (LISA) mission offer a unique way to explore these triple systems (2023LRR....26....2A). By detecting the gravitational wave signals from double white dwarfs, LISA could uncover a population of systems formed through the triple evolution channel, which are inaccessible to electromagnetic observations. This capability has the potential to provide new insights into the formation mechanisms of double white dwarfs, their contribution to Type Ia supernovae (1984ApJS...54..335I; 2024A&A...691A..44K), and the broader implications for the chemical evolution of the Galaxy (1997nceg.book.....P).
@@
LISA is expected to detect $\sim\mathcal{O}(10^{6})$ of Galactic double white dwarfs as part of an unresolved confusion gravitational wave background and individually resolve around $10^{3}$$10^{4}$ of the “loudest” double white dwarfs (e.g., 2017MNRAS.470.1894K; 2019MNRAS.490.5888L; 2021MNRAS.500.4958W; 2023ApJ...945..162T; 2023A&A...669A..82L; 2024arXiv240520484T). While previous studies have focused on double white dwarfs formed from the evolution of isolated binary stars, increasing evidence suggests that hierarchical triple systems, in which a close inner binary is orbited by a distant tertiary companion, may also play a significant role in the formation of double white dwarfs (2020A&A...640A..16T; 2024ApJ...969...68H; 2024arXiv240706257S).
@@
There is mounting observational evidence that stars often form with bound companions, with a binary fraction of $30\,\%$ and a triple fraction of $10\,\%$ for F- and G-type stars (i.e., with masses $\sim 1\,\mathrm{M_{\odot}}$, 2008MNRAS.389..869E; 2010ApJS..190....1R; 2014AJ....147...87T; 2017ApJS..230...15M; 2023ASPC..534..275O. Moreover, the inner binaries in triple systems tend to be in closer orbits than the binaries found in isolated systems, which increases the probability for some binary interactions (2020A&A...640A..16T). In this paper, we show that triple systems offer a greater number of evolutionary pathways for forming short-period inner binaries compared to isolated binary systems. Approximately 10% of white dwarfs are found in binary systems where both components are white dwarfs (i.e., double white dwarfs; 1999MNRAS.307..122M; 2018MNRAS.476.2584M; 2020A&A...638A.131N).
@@
Previous studies have explored the potential for detecting tertiary companions in LISA data (e.g., 2008ApJ...677L..55S; 2018PhRvD..98f4012R; 2019NatAs...3..858T). Similar to electromagnetic observations, the motion of the double white dwarf around the center of mass of the triple system modulates the gravitational wave frequency through the Doppler effect. This modulation produces a periodic shift, causing the observed frequency to oscillate around the intrinsic frequency of the inner binary. Recent studies have focused on leveraging this effect to detect substellar mass tertiaries, such as exoplanets and brown dwarfs (2019NatAs...3..858T; 2019A&A...632A.113D; 2021AJ....162..247K; 2022MNRAS.517..697K). Thus, previous studies have primarily focused on either the isolated binary population of double white dwarfs or the possibility of detecting a third star. Our work is the first evolutionary population synthesis study of Galactic double white dwarfs resulting from triple evolution. In addition, our study assesses the impact of the triple evolution channel to the LISA’s astrophysical noise background, thereby influencing the $\rho$ of all other gravitational wave sources.
@@
We aim to quantify the contribution of the triples to the population of double white dwarfs detectable by LISA. We combine population synthesis models using the MSE code (2021MNRAS.502.4479H) with cosmological simulations from the TNG50 project (2019MNRAS.490.3234N; 2019MNRAS.490.3196P) to construct a representative model of the Galactic double white dwarf population. Our study addresses two critical questions: 1) What fraction of double white dwarfs detectable by LISA originates from the triple evolution channel? 2) Can LISA detect the dynamical effects of the third star in these triple systems?
@@
The paper is structured as follows. In Sect. 2 we explain our methodology. In Sect. 3 we describe the evolutionary pathways of triples that lead to the formation of LISA double white dwarfs. We detail the population properties of LISA double white dwarfs from isolated binaries and triples in Sect. 4, investigating prospects for direct detection of the third star in Sect. 4.4. Finally, we discuss the results in Sect. 5 and summarize our findings in Sect. 6.
@@ 2 Methods
All simulations were performed using the publicly available population synthesis code MSE [^1] https://github.com/hpreece/mse. (2021MNRAS.502.4479H). Our set of simulations of hierarchical triples consists of a main run with a choice of default parameters and three model-variant runs. In each simulated data set, we evolve the triples from the start of the zero-age main-sequence until a maximum integration time $t_{\rm max}=14\,\rm Gyr$. The main triple data set consists of $10^{5}$ systems, where we adopt a common envelope efficiency parameter $\mathrm{\alpha_{CE}=1}$, and all three stars are assumed to have formed at solar metallicity $Z=\rm Z_{\odot}=0.02$. This initial population results in $3\times 10^{3}$ LISA double white dwarfs, with a $\sim 2\%$ Poisson uncertainty in this number. In each of the three model-variant runs, we simulate $10^{4}$ systems and either vary the common envelope efficiency parameter as $\mathrm{\alpha_{CE}=0.1}$ and 10 or change the metallicity of the stars to subsolar $Z=0.1\,\mathrm{Z_{\odot}}$. The effects of the chosen parameters are discussed in Sect. 5. In addition to the triple runs, we simulate a population of $10^{5}$ isolated binaries to compare the impact of tertiary companions. Additionally, we model the inner binaries of all triples without their tertiary stars to assess their influence on the resulting LISA double white dwarf population. We expect that the isolated binary population differs significantly from the inner binary population of triples (2023ApJ...950....9R). The primary distinction is that the inner binaries have much more compact semi-major axes due to the dynamical stability constraints imposed by the tertiary star.
@@
2.1 Multiple stellar evolution code
@@
We used the population synthesis approach using the code MSE to model the stellar evolution, binary interactions (tides, mass transfer), dynamical perturbations from higher-order companions in multiple systems, and flybys from ambient stars. MSE is a C/C++ code with a Python interface that can handle any number of stars as long as they start in a hierarchical arrangement. MSE uses a hybrid approach that switches between the secular approximation for dynamically stable orbits (that satisfy the criterion of 2001MNRAS.321..398M) and direct $N$-body integration using MSTAR (2020MNRAS.492.4131R) for dynamically unstable orbits. Throughout the evolution post-Newtonian terms are included to 2.5 order in the secular approximation and to 3.5 order for the direct $N$-body integration.
@@
To follow the evolution of single stars, MSE relies on the fitting formulae from 1996MNRAS.281..257T; 2000MNRAS.315..543H, while binary interactions such as tides, wind mass transfer, stable mass transfer episodes, and common envelope (CE) evolutions are computed using modified prescriptions of 2002MNRAS.329..897H. We briefly explain the physical handling of these key processes below. For more detailed explanations see 2021MNRAS.502.4479H.
@@
Stable mass transfer: In MSE, mass transfer stability is determined either by the critical mass ratio criterion or by comparing mass transfer and dynamical timescales. The critical mass ratio depends on the donor’s stellar type (2021MNRAS.502.4479H). We assume fully conservative mass transfer ($\beta_{\mathrm{MT}}=1$), meaning no mass is lost from the system. While MSE generally follows 2002MNRAS.329..897H for binary interactions, it differs in treating mass transfer in eccentric orbits. 2002MNRAS.329..897H assumes tides are always efficient in circularizing the orbit. However, in triple systems, the eccentricities are excited secularly. We follow the analytical model from 2019ApJ...872..119H to model mass transfer at periastron in eccentric orbits
@@
Common envelope evolution: Unstable mass transfer/CE evolution in MSE follows the $\alpha_{\mathrm{CE}}$ prescription (1976IAUS...73...75P; 1976IAUS...73...35V; 1988ApJ...329..764L; 1993PASP..105.1373I; 2002MNRAS.329..897H). The code solves for orbital energies before and after the CE phase, parameterizing the envelope ejection efficiency by $\alpha_{\mathrm{CE}}$ and adopting a binding energy factor $\lambda$ for each donor star taken from fits described in 2014A&A...563A..83C, assuming all the ionization energy stored in each envelope is useful in unbinding that envelope. For the main runs, we adopt $\alpha_{\mathrm{CE}}=1$. The post-CE semi-major axis is determined from the corresponding orbital energy. The timescale of mass ejection from a CE in the inner binary is relevant for the evolution of the outer orbit, and we assume that ejected common-envelope material leaves the system in a timescale of $10^{3}\,\mathrm{yr}$.
@@
Merger or collision: A “failed” CE can result in the merger of two stars. This occurs if the post-CE semi-major axis is too small for either star to avoid Roche lobe overflow. Beyond post-CE mergers, MSE also accounts for physical collisions when the sum of the stellar radii exceeds the semi-major axis or when a periapsis collision occurs in an eccentric orbit. The properties of the merger remnant are assigned following 2002MNRAS.329..897H.
@@
Triple common envelope (TCE) evolution: Mass transfer from a third star onto the inner binary can lead to a TCE. If CE conditions are met for the third star, MSE employs “circumstellar triple CE evolution,” allowing the third star to fill its Roche lobe around the inner binary and undergo unstable mass transfer. This follows a similar approach to the prescription proposed by 2020MNRAS.498.2957C. The final outer semi-major axis is estimated using an $\alpha_{\mathrm{CE}}$ prescription, assuming the inner binary remains intact and does not change. However, CE modeling is uncertain (see 2013A&ARv..21...59I for a review), and TCE evolution is even more so, requiring cautious interpretation of results. For hydrodynamical simulations of triple CE outcomes, see, for example, 10.1093/mnras/staa3242.
Flybys: MSE also includes the gravitational perturbations from stellar flybys in the vicinity of the system using the impulsive approximation. The flyby mass is sampled from a Kroupa initial mass function (2001MNRAS.322..231K), and encounters are randomly sampled within an encounter sphere of radius $R_{\text{enc}}=10^{5}\,au$ with velocities drawn from a Maxwellian velocity distribution with dispersion $\sigma_{\star}=30$ km s-1. The number density of flybys accounts for the low-density environments, assuming $n_{\star}=0.1\,\mathrm{pc}^{-3}$. These flybys become significant for the evolution of the system if the semi-major axis of the outer orbit exceeds about $10^{3}\,\rm au$ (2010MNRAS.401..977J; GrishinPerets; Stegmann:2024rnk).